OpenStudy (anonymous):

Which of the following circle equations represents a circle whose center is at (0, 0)?Select one of the options below as your answer: A. x2 + y2 = 45 B. (x − 3)2 + (y − 5)2 = 0 C. (x + 3)2 + (y − 3)2 = 0 D. x2 + (y − 1)2 = 1

OpenStudy (whpalmer4):

Use the formula I gave you on the last problem. Once again, the formula for a circle with radius \(r\) and center at \((h,k)\) is\[(x-h)^2+(y-k)^2=r^2\] Your center is at \((0,0)\) so \(h=0\) and \(k=0\). Substitute those values, simplify the equation and compare with the answer choices.

OpenStudy (anonymous):

so its c?

OpenStudy (whpalmer4):

Okay, let's do what I just suggested and rewrite the formula with \(h=0\) and \(k=0\): \[(x-0)^2+(y-0)^2=r^2\]\[x^2+y^2=r^2\] Here are the answer choices: A: \(x^2+y^2=45\) B: \((x − 3)^2 + (y − 5)^2 = 0\) C. \((x + 3)^2 + (y − 3)^2 = 0\) D. \(x^2 + (y − 1)^2 = 1\) Which of those choices most resembles \(x^2+y^2=r^2\)?

OpenStudy (anonymous):

ohhhhh A

OpenStudy (whpalmer4):

Yes! :-)

OpenStudy (anonymous):

yay thanks :D Im stressin i have 20 minutes left and 7 more questions

OpenStudy (whpalmer4):

hopefully you've just figured out how to do them :-)

OpenStudy (anonymous):

thanks